OpenStudy (anonymous):

Find the average rate of change of the function: f(x)=0.5x-x^2 from x=-2 to x=4

OpenStudy (anonymous):

so the y's or the f(x)'s are equal to the x's?

OpenStudy (anonymous):

was just trying to figure out how to plug it in correctly, I remembered the formula. But didn't recall how to get the f(x)'s.

OpenStudy (lalaly):

\[f(4)=0.5(4)-4^2\]

OpenStudy (anonymous):

nice thanks. so now I figure the equation using the outcome right?

OpenStudy (anonymous):

f(4) = -14

OpenStudy (anonymous):

doesn't "rate" mean to differentiate? i think you need to find dy/dx and then put in the value at those points. No?

OpenStudy (anonymous):

so I would input -14 in to the \[f(x _{2}) = -14\] So \[\frac{-14-3}{4--2}\] so the average rate of change is: \[\frac{-17}{6}\]

OpenStudy (anonymous):

Is that correct or no?

OpenStudy (anonymous):

Can't be right suppose to equal -3/2 but I can't figure out how to get that answer.

OpenStudy (anonymous):

dy/dx= 0.5-2x, nowput x= -2 and x= 4, to get gradient which is 4.5 and 7.5, take its average i.e 12/2 = 6

OpenStudy (anonymous):

aroc or average rate of change = slope which is y2-y1/x2-x1 which will give your your average rate of change.

OpenStudy (anonymous):

which in turn = f(x2)-f(x1)/x2-x1

OpenStudy (anonymous):

oh okay